OpenStudy (anonymous):

How much would you have to invest at 8.5 percent interest,compounded quarterly, to have 5,000 after 6 years?

jimthompson5910 (jim_thompson5910):

Hint: Use the formula A = P(1 + r/n)^(nt) where A = final amount P = amount invested r = interest rate n = number of times compounded per year t = number of years

jimthompson5910 (jim_thompson5910):

In this case, A = 5000 P = x (this is unknown) r = 0.085 n = 4 t = 6 Plug all this in and solve for x to find the amount to invest

OpenStudy (campbell_st):

is you interest rate per annum or per quarter..?

OpenStudy (anonymous):

im guessing quarterly,,,im not sure....that's the question word for word.

jimthompson5910 (jim_thompson5910):

it's most likely the annual interest rate

OpenStudy (campbell_st):

well I'd use the method shown above... lol... but it should be stated

OpenStudy (anonymous):

i've already tried the above method, and i failed horribly lol

jimthompson5910 (jim_thompson5910):

A = P(1 + r/n)^(nt) 5000 = P(1 + 0.085/4)^(4*6) 5000 = P(1 + 0.02125)^(24) 5000 = P(1 + 0.02125)^(24) 5000 = P(1.02125)^(24) 5000 = P(1.65641696) 5000/(1.65641696) = P 3018.5636 = P P = 3018.5636 P = 3018.56 So you need to invest 3018.56

OpenStudy (anonymous):

thank you so much love <3

jimthompson5910 (jim_thompson5910):

you're welcome, I'm glad I could help you out

jimthompson5910 (jim_thompson5910):

and hopefully it made sense?

OpenStudy (anonymous):

yes, it did, i wasnt sure how to solve it, i kept getting some weird numbers :P

jimthompson5910 (jim_thompson5910):

alright that's great